Find the future value of an initial investment of $5,000 invested at 7% compounded annually, for 25 years.
the formula is: FV=PV(1+i)^n when
FVn is the Futer value of the investment after n years
Pv is the initial amount of the investment
i is the interest rate.
n is the length of the investment, in years.
I have set the equation up like this
FVn= 5,000(1+7%)^25
FVn=5,000(1+0.07)^25
FVn=5,000(1.07)^25
Is this the correct set up?
Now I know since I have the 25 exponent, I have to multiply the 1.07 times itself 25 times, can someone help me from that step. I do not have a scientific calculator so I really need assistance with the multiplication of the exponent and the solution to the problem.
thank you in advance for your assistance

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

FVn=5,000(1.07)^25
Is this the correct set up?
Now I know since I have the 25 exponent, I have to multiply the 1.07 times itself 25 times, can someone help me from that step. I do not have a scientific calculator so I really need assistance with the multiplication of the exponent and the solution to the problem.
--------------
I think you can find a scientific calculator online;
use Google to find one.
Here's one really great site: http://my.hrw.com/math06_07/nsmedia/tools/Sci_Calculator/Sci_Calculator.html
--------------
Your work is correct.
FVn = 5000(1.07)^25
FVn = 5000*5.4274
FVn = 27137.16
====================
Cheers,
Stan H.

Answer by mananth(16949)

Principal P = 5000
Amount= A
years=n 25
compounded 1 times a year
Rate = 7.00 0.07
Amount = P*((n+r)/n)^n

Amount =5000*(1+0.07)^25*1
Amount =5000*(1 +0.07)^ 25
5000 *(1.07)^25
Amount = 27137.16